How many 13-card hands having exactly 11 diamonds can be dealt?
Answer by stanbon(75887)

# of ways to pick 11 diamonds: 13C11 = 13*12/1*2 = 78
---------------
# of ways to pick 2 other cards: 39C2 = 39*38/1*2 = 19*39 = 741
----------------
# of ways to have a 13-card hand with exactly 11 diamonds: 78*741 = 57,798
=====================
Cheers,
Stan H.